# Who You Are Matters: Bridging Topics and Social Roles via LLM-Enhanced Logical Recommendation

### v2

@@ 2.3.2 Tag-based Learning Augmentation
Tag-Logic Exploration: For the settings of $\mathcal{T}_{u}^{+}$, $\mathcal{T}_{u}^{-}$, and $\mathcal{T}_{i}^{+}$, we offer two alternatives that either use the original tag sets $\mathcal{T}_{u}(0)=\{t|t\in\arg_{t}\text{top-}k[P(t|u)]\}$ (or $\mathcal{T}_{i}(0)=\{t|t\in\arg_{t}\text{top-}k[P(t|i)]\}$ for $\mathcal{T}_{i}^{+}$) that address the recommendation utility (denoted as TagCF-util) or use the extended tag sets $\mathcal{T}_{u}(1)$ (or $\mathcal{T}_{i}(1)$) inferred by the logic graphs that address the interest exploration (denoted as TagCF-expl). For instance, we have a target item that has an initial tag $t=$ “Symphonist” which is logically related to the topic $c=$ “Music Theory” according to $\mathcal{G}^{\text{U2I}\ast}$. Then using $\mathcal{G}^{\text{I2U}\ast}$, we might explore and find that there exists a logic of “Music Theory” $\rightarrow$ “Teacher”, where “Teacher” becomes the extended tag of the item. We provide a detailed description of the general procedure in Appendix B.6 and the confirmatory case study in Appendix D.3.
@@ 3.1.1 Workflow Specification
We conduct an online A/B test on a real-world industrial video recommendation platform to evaluate the effectiveness of TagCF-ut. The platform serves videos for over half a billion users daily, and the item pool contains tens of millions of videos. Figure 3 provides a detailed overview of the implementation of our online recommendation workflow. The tag extraction module, the collaborative logic reasoning module, and the training of all augmented models are offline procedures executed on a daily basis. In contrast, the inference part of the tag-logic integration module is deployed in the last ranking stage (that chooses top-6 scored items as recommendation from 120 candidates from previous stage) for real-time recommendation requests, with preprocessed tag and logic information retrieved from the latest knowledge base. As we have described in Section 2.1 and Section 2.2, we use the cover set solution to achieve stable and efficient inference that fulfills the industrial demand.
@@ 3.2.1 Experimental Setup
Datasets: To further investigate the design choices of TagCF, we include two public datasets mcauley2015image , Books and Movies, as well as an offline dataset from our real-world industrial video sharing platform (i.e., Industry). More details about the datasets and preprocessing can be seen in Appendix C.2.
Evaluation Protocol: We include common ranking accuracy indicators such as $N$ and MRR@$N$, as well as diversity metrics like ItemCoverage@$N$ and GiniIndex@$N$ (denoted as Cover@$N$ and Gini@$N$, respectively). In this paper, we observe $N\in\{10,20\}$. For each experiment across all models, we run training and evaluation for five rounds with different random seeds and report the average performance.
Baselines: We include BPR rendle2012bpr as the standard collaborative filtering method, and include several representative sequential models, namely GRU4Rec hidasi2015session , Bert4Rec sun2019bert4rec , SASRec kang2018self , LRURec yue2024linear , Mamba4Rec liu2024mamba4rec . We also compare with LLM-enhanced recommender approaches: RLM ren2024representation , SAID hu2024enhancing and GENRE liu2024once . See more baseline details in Appendix C.
@@
Dataset Method NDCG@10 NDCG@20 MRR@10 MRR@20 Cover@10 Cover@20 Gini@10$\downarrow$ Gini@20$\downarrow$
@@ 3.2.3 Transferability Test
To validate that the extracted tags (i.e., $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$) and logic graphs (i.e.,$\mathcal{G}^{\text{U2I}\ast}$ and $\mathcal{G}^{\text{I2U}\ast}$) in our industrial solution encapsulate general knowledge, we conduct a cross-task transfer experiment. Specifically, we use the same tag extraction module in Eq.(1) to generate the data-specific tags for Books and Movies data. Then we use the semantic embedding multim3 of these tags to find the closest tags in $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$ so that the tag space is completely aligned. This means that the TagCF solutions for these two public datasets can skip the collaborative logic reasoning module and directly use $\mathcal{G}^{\text{U2I}\ast}$ and $\mathcal{G}^{\text{I2U}\ast}$ for tag exploration. The experimental results are presented in Table 2 and we observe that TagCF variants consistently demonstrate superior recommendation accuracy and diversity in Books and Movies, proving its transferability to other tasks. Note that other LLM-based baselines also use the extracted tag-logic information to enhance recommendation in our experiments, which indicates that the knowledge is transferable to other methods as well.
@@ 3.2.4 Ablation Study
  • Integration methods: To evaluate the individual impact of the three main components in the integration framework (section 2.3), we compare the full TagCF with three alternatives each disables one component in {tag-based encoder, tag-based learning augmentation, and tag-logic inference}, denoted as w/o TE, w/o TA, and w/o TLI, respectively. We show the results on the Industrial dataset in Figure 4, which verifies that all three components contribute to the recommendation accuracy. The same conclusion applies to diversity metrics as well and the results are illustrated in Figure 8 of Appendix C.2. The performance degradation observed in each ablated variant underscores the complementary value of each module within the integrated framework.
@@
  • Effect of $k$: We conduct experiments with a different number of tags extracted for each item ($k\in\{20,50,100,200,\text{full}\}$) and present the results in Figure 9 in Appendix C.2.
@@
In Table 2, we find that TagCF-ut tends to yield greater improvements in accuracy metrics, indicating that the user tag set is potentially more effective and stable in capturing preferences and personalities of users, solving the role identification task. This phenomenon might be related to the fact that user role tags are likely to be stable concepts with better expressiveness, which can be partially explained by the smaller cover set size compared with item tags shown in Table 6 of Appendix D.1. In contrast, item tags may have a shorter lifespan (e.g., a special topic in recent news) and may frequently update even in the cover set. This may also explain the optimal diversity performance of TagCF-it, since the more fine-grained item tag set can contribute more diverse options during training and inference.
@@ 4 Conclusion
In this work, we emphasize the importance of the modeling of users’ roles and the user-item behavior logic in the semantic tag space, and propose a new recommendation paradigm, TagCF, that can effectively extract item/user tags from items with MLLM, infer realistic behavioral logic of users with LLM, and enhance recommendation performance with the tag-logic knowledge. We provide technical details of our efficient and effective solution of TagCF, which has been successfully deployed in our industrial video-sharing platform. We also verified that the extracted knowledge of the logic graph is a general transferable asset to other recommendation tasks and LLM-based augmentation methods. Compared to the item tag set, the user role tags are empirically more stable and more expressive in representing personalities, improving recommendation accuracy, indicating a statistically better design choice than the traditional item-tag-based methodology. For the future of the recommender system, this work opens a new direction that aims to achieve explicit modeling of the user’s social role and behavioral logic in real life.
Appendix A Related Work
A.1 Collaborative Filtering
Collaborative filtering (CF), one of the most successful recommendation approaches, continues to attract interest in both academia and industry. Over time, CF has evolved from traditional methods [44, 33, 7, 5, 39] to advanced techniques incorporating sequences [25, 29, 48] and graph structures [51, 24]. Traditional CF methods, such as matrix factorization techniques [5] and matrix factorization (MF) [39], have been effective at learning latent user and item representations. Sequential CF extends this by modeling the temporal order of user interactions, such as GRU4Rec [25] employs Recurrent Neural Networks (RNNs) to capture temporal dependencies, while SASRec [29] introduces self-attention to focus on historical interactions. BERT4Rec [48] uses bidirectional Transformers to consider both past and future contexts. Graph-based CF methods like NGCF [51] and LightGCN [24] have also gained attention in recent years. Besides, self-supervised learning approaches [58, 9, 60] have been explored to enhance CF by learning robust representations. However, these methods often ignore user roles and logical relationships between characteristics.
Meanwhile, some personality-aware filtering methods incorporate user traits through neighborhood filtering [30, 16, 15] or matrix factorization extensions [30, 16]. However, due to the popularity of Big-Five model in the literature of psychology [21], the majority of the works used Big-Five personality model to represent the user’s personality, while the choice of the most suitable personality definition that satisfy the requirements of the recommendation application still needs further investigation. Recent works [32, 57, 45] have attempted to leverage LLMs for personalized recommendations and user interest interpretation. While progress has been made, existing approaches still overlook explicit modeling of user roles and their logical relationships. In this work, we aim to address these gaps by bridging topics and social roles via LLMs-enhanced logical recommendation within the CF framework.
A.2 LLM-based Recommendation
LLM-enhanced Recommender. Many current works have explored how to apply the LLM to generate auxiliary knowledge for enhancing traditional RS. LLMRG [52] fabricates prompts to construct chained graph reasoning from LLM to augment the recommendation model. LLMHG [12] first leverages LLMs to deduce Interest Angles (IAs) and categorize movies into multiple categories within each IA to construct a multi-view hypergraph. SAGCN  [36] uses a chain-based prompting strategy to extract semantic interactions from LLM for each review and introduces a semantic aspect-based graph convolution network to enhance the user and item representations by leveraging these semantic aspect-aware interactions. LLM-KERec [62] uses LLM to identify the complementary relationships of an item knowledge graph. Subsequently, they train an entity-entity-item weight decision model which is then used to inject knowledge into the ranking model by using the real exposure and click feedback of complementary items. Nevertheless, current methods of using LLM for data enhancement primarily focus on the meta-features, neglecting knowledge from the user side and the logic rationale between user-item interactions. This limitation hinders their ability to facilitate traditional recommenders to capture semantic and representative collaborative information.
LLM as Recommender itself. Recently, LLMs have demonstrated remarkable performance across a wide range of recommendation tasks. P5 [20] and M6Rec [13] finetune LLM by modeling recommendation tasks as natural language processing tasks. ChatRec [17] employs LLMs as a recommender interface for conversational multi-round recommendations. TALLRec [6] designs a customized parameter-efficient tuning process for recommendation tasks on LLM with a LoRA architecture. HLLM  [11] uses an item LLM to encode text features, feeding its embeddings to a user LLM for recommendations. Compared to LLM-enhanced recommender, this paradigm’s computational cost (both for training and inference) is much higher and the industry-deployable solution is still an open question [14]. As we will discuss in section 2.3.1, this research direction focuses on the improvement of sequential models, which is complementary to our proposed knowledge extraction and augmentation framework.
Appendix B TagCF Specification
B.1 Task Formulation

### v3

@@ 3.3.2 Tag-based Learning Augmentation
Tag-Logic Exploration: For the settings of $\mathcal{T}_{u}^{+}$, $\mathcal{T}_{u}^{-}$, and $\mathcal{T}_{i}^{+}$, we offer two alternatives that either use the original tag sets $\mathcal{T}_{u}(0)=\{t|t\in\arg_{t}\text{top-}k[P(t|u)]\}$ (or $\mathcal{T}_{i}(0)=\{t|t\in\arg_{t}\text{top-}k[P(t|i)]\}$ for $\mathcal{T}_{i}^{+}$) that address the recommendation utility (denoted as TagCF-util) or use the extended tag sets $\mathcal{T}_{u}(1)$ (or $\mathcal{T}_{i}(1)$) inferred by the logic graphs that address the interest exploration (denoted as TagCF-expl). For instance, we have a target item that has an initial tag $t=$ “Symphonist” which is logically related to the topic $c=$ “Music Theory” according to $\mathcal{G}^{\text{U2I}\ast}$. Then using $\mathcal{G}^{\text{I2U}\ast}$, we might explore and find that there exists a logic of “Music Theory” $\rightarrow$ “Teacher”, where “Teacher” becomes the extended tag of the item. We provide a detailed description of the general procedure in Appendix A.6 and the confirmatory case study in Appendix C.3.
@@ 4.1.1 Workflow Specification
We conduct an online A/B test on a real-world industrial video recommendation platform to evaluate the effectiveness of TagCF-ut. The platform serves videos for over half a billion users daily, and the item pool contains tens of millions of videos. Figure 3 provides a detailed overview of the implementation of our online recommendation workflow. The tag extraction module, the collaborative logic reasoning module, and the training of all augmented models are offline procedures executed on a daily basis. In contrast, the inference part of the tag-logic integration module is deployed in the last ranking stage (which chooses top-6 scored items as recommendation from 120 candidates from the previous stage) for real-time recommendation requests, with preprocessed tag and logic information retrieved from the latest knowledge base. As we have described in Section 3.1 and Section 3.2, we use the cover set solution to achieve stable and efficient inference that fulfills the industrial demand.
@@ 4.2.1 Experimental Setup
Datasets: To further investigate the design choices of TagCF, we include two public datasets ni2019justifying , Books and Movies, as well as an offline dataset from our real-world industrial video sharing platform (i.e., Industry). More details about the datasets and preprocessing can be seen in Appendix B.2.
Evaluation Protocol: We include common ranking accuracy indicators such as NDCG@$N$ and MRR@$N$, as well as diversity metrics like ItemCoverage@$N$ and GiniIndex@$N$ (denoted as Cover@$N$ and Gini@$N$, respectively). In this paper, we observe $N\in\{10,20\}$. For each experiment across all models, we run training and evaluation for five rounds with different random seeds and report the average performance.
Baselines: We include BPR rendle2012bpr as the standard collaborative filtering method, and include several representative sequential models, namely GRU4Rec hidasi2015session , Bert4Rec sun2019bert4rec , SASRec kang2018self , LRURec yue2024linear , Mamba4Rec liu2024mamba4rec . We also compare with LLM-enhanced recommender approaches: RLM ren2024representation , SAID hu2024enhancing and GENRE liu2024once . See more baseline details in Appendix B. We follow RecBole zhao2021recbole as the implementation backbone and reproduce all baselines with hyper-parameters from either the original setting provided by authors or fine-tuning using validation.
@@
[…]
@@ 4.2.3 Transferability Test
[…]
@@ 4.2.4 Ablation Study
  • Integration methods: To evaluate the individual impact of the three main components in the integration framework (section 3.3), we compare the full TagCF with three alternatives each disables one component in {tag-based encoder, tag-based learning augmentation, and tag-logic inference}, denoted as w/o TE, w/o TA, and w/o TLI, respectively. We show the results on the Industrial dataset in Figure 4, which verifies that all three components contribute to the recommendation accuracy. The same conclusion applies to diversity metrics as well and the results are illustrated in Figure 8 of Appendix B.2. The performance degradation observed in each ablated variant underscores the complementary value of each module within the integrated framework.
@@
  • Effect of $k$: We conduct experiments with a different number of tags extracted for each item ($k\in\{20,50,100,200,\text{full}\}$) and present the results in Figure 9 in Appendix B.2.
@@
In Table 2, we find that TagCF-ut tends to yield greater improvements in accuracy metrics, indicating that the user tag set is potentially more effective and stable in capturing preferences and personalities of users, solving the role identification task. This phenomenon might be related to the fact that user role tags are likely to be stable concepts with better expressiveness, which can be partially explained by the smaller cover set size compared with item tags shown in Table 6 of Appendix C.1. In contrast, item tags may have a shorter lifespan (e.g., a special topic in recent news) and may frequently update even in the cover set. This may also explain the optimal diversity performance of TagCF-it, since the more fine-grained item tag set can contribute more diverse options during training and inference.
@@ 5 Conclusion
In this work, we emphasize the importance of the modeling of users’ roles and the user-item behavior logic in the semantic tag space, and propose a new recommendation paradigm, TagCF, that can effectively extract item/user tags from items with MLLM, infer realistic behavioral logic of users with LLM, and enhance recommendation performance with the tag-logic knowledge. We provide technical details of our efficient and effective solution of TagCF, which has been successfully deployed in our industrial video-sharing platform. We also verify that the extracted knowledge of the logic graph is a general transferable asset to other recommendation tasks and LLM-based augmentation methods. Compared to the item tag set, the user role tags are empirically more stable and have more potential in improving recommendation accuracy, shedding light on an alternative design choice to the traditional item-tag-based methodology, posing new challenges to recommender systems.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in the abstract and introduction accurately reflect the contributions and scope of our paper. The abstract succinctly summarizes our approach, the tag-based logic filtering (TagCF) framework. It highlights the importance of user social role modeling and the integration with tag-based encoder, tag-based learning augmentation, and the tag-logic inference extension, providing a clear overview of the method’s novelty and effectiveness. The introduction elaborates on the motivation, background, and significance of our contributions, ensuring that the claims align with the detailed discussions and results presented in the subsequent sections of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations and future directions in the Appendix F.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: we provide the full set of assumptions and a complete (and correct) proof in our methodology part 3. All the theorems, formulas, and proofs in the paper are numbered and cross-referenced.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper describes the architecture clearly. We fully disclose all the information needed in Appendix A and Appendix B. This includes details about the datasets, experimental setups, hyperparameters, evaluation metrics, and model specification.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have prepared both the dataset and source code, and will release them promptly upon paper publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental settings are presented in the core of the paper. And full details are provided appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have statistical significance in the main experiment.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: : We indicate the sufficient information on the type of GPU compute workers, memory and time of execution.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We checked and ensured that our paper conforms with the NeurlPS Code of Ethics in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our paper discusses the potential positive and negative societal impacts of our work in the "Broader Impacts" section E. By acknowledging these impacts, we provide a balanced view of our work and suggest mitigation strategies to address potential negative outcomes.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creator or original owner of the assets (e.g., code, data, models) used in the paper is properly credited, and the license and terms of use are explicitly mentioned and appropriately respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We are not involved in these risks as we are only engaged in recommendation tasks. Our research does not involve human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our research employs LLMs as an integral and novel component of the core methodology, specifically, we utilize LLMs to extract interpretable user tags and item tags from user behavior data, and we employ LLMs to generate the tag-based logic graph through chain-of-thought reasoning. We have documented the LLM usage details in Section 3.1 and Section 3.2.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A TagCF Specification
A.1 Task Formulation